# ESGSenticNet: A Neurosymbolic Knowledge Base for Corporate Sustainability Analysis

### v1

@@ 4.1 Methodology
Sustainability disclosures are tokenised into separate sentences, with each sentence transformed into embeddings by Sentence-BERT (S-BERT) [45]. BERTopic is then run on these embeddings, with the configuration of KeyBERT [20]. The topic words are then provided to sustainability experts (A) to facilitate their development of the taxonomy for ESGSenticNet (table 1). Examples of topic words and their inferred topics can be found in table 1. The result is a taxonomy (table 2) with the following features.
  • Realistic: Given that taxonomy development leverages empirical insights from the sustainability reports themselves, categories can capture content within sustainability reports that might otherwise be overlooked by standard frameworks. For instance, ‘Green Technology’ is derived as a taxonomy category, while not being explicitly emphasised in standard reporting frameworks (i.e. GRI [33]).
  • Activities Cross-Theme: To capture the interconnectedness of sustainability dimensions [38], categories include areas of a firm’s activities (‘operations’, ‘development’ etc.), on top of sustainability impacts (‘emissions control’, ‘waste management’ etc.). In contrast with existing frameworks, this offers insight into the sustainability of specific business activities.
  • Granularity: The taxonomy has a hierarchical nature to enable analysis at different levels of granularity.
Topic keywords Category
@@
Table 1: Examples of Categories Inferred from Topic Keywords
4.2 Resultant Taxonomy
Table 2 shows the hierarchical taxonomy derived for ESGSenticNet. It contains the definitions of the pillars (Environmental, Social, Governance) and topics used in ESGSenticNet. Along with the pillars, the broad topics, and cross-broad topics are bolded, with their constituent sub-topics highlighted in bullet points. Beside the broad and cross-broad topics, a letter denotes which pillar the topic falls under – (E) Environmental, (S) Social, (G) Governance.
Category Description
Table 2: ESGSenticNet Sustainability Taxonomy
5 ESGSenticNet
ESGSenticNet comprises knowledge triplets with the format of (concept, relation, category), following from [27]. To form the knowledge triplets, concepts within our knowledge base are labelled according to their relations (table 3) with the categories in our taxonomy (table 2). The relations, ‘supports’ and ‘undermines’, clarify how concepts advance or impede crucial aspects of corporate sustainability. In contrast, the relation ‘aligns with’ does not consider a concept’s impact, highlighting a concept’s general connection with broader pillars of environmental, social or governance. Additionally, categories are divided according to their different category types–pillar, broad, cross-broad, sub, cross-sub, as shown in table 4, with relations assigned according to rules (section 5.1).
ESGSenticNet is constructed through a pipeline that involves three main phases: parsing concepts from the text corpus (section 5.2), processing parsed concepts to ensure coherence (section  5.3), labelling processed concepts according to their relation with respect to our developed taxonomy (sections 5.55.65.45.7). Concept parsing involves extracting candidate concepts from sustainability disclosures through dependency parsing and part-of-speech tagging. From these candidates, coherent concepts are filtered through text parsing followed by leveraging GPT-4 [3] (section 5.2). The filtered concepts undergo a labelling methodology that we have developed to minimise computational cost while enhancing the diversity of labelled concepts. Accordingly, this labelling methodology (figure 1 1 through 4) entails leveraging semantic graphs, clustering techniques, seed selection algorithms, and label propagation. This methodology allows us to label a large amount of diverse concepts despite performing inference on a subset of the concept samples. As a result, the extensiveness and diversity of our knowledge base is significantly increased, which is imperative to its application for corporate sustainability analysis.
@@ 5.1 ESGSenticNet Relations & Rules
supports The concept directly advances the sustainability topic with a clear and immediate positive impact.
undermines The concept directly impedes the sustainability topic with a clear and immediate negative impact.
@@
Category Type Categories
Pillar ‘Environmental’, ‘Social’, ‘Governance’.
Sub Children categories of broad categories
Cross-Sub Children categories of cross-broad categories
Table 4: Categories and category types
ESGSenticNet is defined with the following relations (table 3), category types (table 4), as well as the following rules.
  • Pillar Assignment: Each concept ‘aligns with’ a specific pillar, mapping the concepts to the broader sustainability context. Single Label within Category Types: Each concept can hold only one relation with one category within each category type.
  • Cross-Labels between Parent & Children: Each concept can hold relations with higher-level categories (‘broad’ and ‘cross-broad’), as well as with more detailed categories (‘sub’ and ‘cross-sub’).
  • Cross-Labels between Cross & Non-Cross: Concepts that ‘align with’ the ‘environmental’ pillar can hold relations with ‘cross-broad’ and ‘cross-sub’ categories, in addition to ‘broad’ and ‘sub’ categories. This reflects the multi-faceted nature of environmental concepts, which can simultaneously impact environmental stewardship (represented by ‘cross-broad’, ‘cross-sub’ categories) and strategic business activities (represented by ‘broad’ and ‘sub’ categories).
@@ 5.2 ESG Concept Parser
Multiword expressions are linguistic phenomena that refer to common linguistic constructions such as idioms, collocations and sayings [16]. In the context of sustainability analysis, multiword expressions frequently communicate key sustainability ideas [51]. We refer to these multiword expressions as sustainability ”concepts”. Extracting these concepts or ”concept extraction” has been explored for metaphors [36], finance [18] and sentiment analysis [10]. Yet, to the best of our knowledge, concept extraction within the sustainability domain has not been studied. Therefore, to capture these sustainability concepts, a new concept parser is required, with an explicit focus on the distinct elements of sustainability discourse. These include the following designs:
@@
To effectively extract these unique aspects of sustainability language, we formalise sustainability concepts, or ESG concepts, through the following syntactic structure–verbs acting on complex noun phrases, where the words involved function as modifiers, descriptive and prescriptive elements. Specifically, the concepts are two or three word phrases, typically beginning with a $verb$, followed by $noun+noun$, or $adjective+noun$ combinations with specific grammatical dependencies. These dependencies include: $verb-noun$ connection via (nsubj, obj, obl), $adjective-noun$ or $noun-noun$ connection via (compound, amod, nn, appos, flat, nmod). This framework underpins the ESG Concept Parser algorithm (1), designed to capture key ideas within sustainability disclosures, particularly sustainability efforts. For instance, ‘improve (VB) workplace (NN) safety (NN)’, where ‘workplace’ is the $compound$ of ‘safety’ and ‘safety’ is the $dobj$ of ‘improve’.
@@
1:Input: Sentence from sustainability disclosure
@@
16:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+noun+noun combination
@@
19:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+adj+noun combination
@@
5.3 Processing concepts for coherence
To minimise computational cost, parsed ESG concepts from algorithm (1) are filtered to retain the top 110k most frequently occurring ones. Of these, we eventually annotate 23k concepts based on our estimated computational resources, which dictates the number of seeds we select within our semantic graph. Prior to the annotation process, these concepts must undergo processing to ensure their coherence. This involves utilising GPT-4o in few-shot setting (Brown, 2020) to ensure their coherence. GPT-4o is leveraged due to the capacity of large generative LLMs to infer [58], as well as their powerful natural language understanding capabilities [62]. Using GPT-4o to process concepts for coherence involves conducting the following inference tasks in order: (1) reorder words, if required, to improve the coherence of the ESG concept, (2) determine if the subsequent ESG concept is intelligible. We provide the prompt for (1) and (2) through figures 3 and 4 respectively.
@@ 5.3 Processing concepts for coherence
5.4 Semantic Graph Construction
@@
Processed ESG concepts undergo labelling to derive their relationships with different sustainability topics. Given the exceedingly large number of ESG concepts, annotating all concepts would be prohibitively expensive. To overcome this, we devise a semi-supervised framework that enables a large volume of diverse ESG concepts to be labelled from annotating only a subset of all ESG concepts. Via a semantic graph, we leverage the semantic similarities between ESG concepts to propagate labels from annotated concepts (seeds) to unlabelled concepts (non-seeds). This graph is constructed by transforming each concept $cp_{i}$ into embeddings $e_{i}$ through S-BERT [45], with edges constructed by leveraging a cosine similarity threshold of 0.80 between each node $e_{i}$.
5.5 Seed Selection
Within our semantic graph, seeds are selected for annotation through data clustering and graph analysis. Given the semi-supervised label propagation from seeds to non-seeds, we formulate the following objectives for seed selection: (I) Seeds must be semantically diverse to ensure that diverse labelled samples are yielded post label propagation. This enhances the variety and comprehensiveness of labelled ESG concepts in our knowledge base. (II) Seeds must propagate their labels to as many non-seeds as possible. This maximises the number of labelled samples we can attain while minimising annotation costs.
5.5.1 Data Clustering
@@
To accomplish (I), the embeddings of ESG concepts $e_{i}$ are clustered to enable seed selection from diverse semantic spaces. This clustering involves projecting the embeddings to a lower dimensional space via UMAP [37], before running HDBSCAN [12] on these projected embeddings. While HDBSCAN is flexible to the shape and number of clusters, and robust to noise [12], reducing the dimensionality of embeddings ensures good clustering performance. This is because HDBSCAN relies on distance-based metrics which lack meaningfulness in the high dimensional spaces of textual embeddings [4]. The embeddings are projected to lower dimensions for the sole purpose of clustering. In our later section, label propagation leverages the default dimensions of S-BERT embeddings.
Clusters are evaluated via a Confidence Quality Index ($CQI$), which denotes the proportion of embeddings, $e$, out of all embeddings, $V$, with a clustering confidence level $\alpha_{e}$ beyond 0.60. For our use of HDBSCAN, this simple metric is more appropriate than others like Slihouette Score [47], which assumes that all embeddings belong to a cluster. Hyperparameter tuning is done via random search, with the best run having a $CQI$ of 0.66, UMAP parameters $n$_$neighbours$$=$32, $n$_$components$$=$2, and the HDBSCAN parameter $min$_$cluster$_$size$$=$2.
5.5.2 Graph Analysis
@@
To accomplish (II), graph analysis is utilised to select seeds that are the most consequential nodes. Given that edges facilitate label propagation, a node is considered more consequential if it possesses a greater number of unique edges. Unique edges are defined as connections to nodes that do not share an edge with the already selected seed nodes within the same cluster. We formulate the local consequential score $Q(e_{i})$, where $S_{c_{k}}$ is the set of nodes already selected as seeds within the same cluster, $c_{k}$, as the node $e_{i}$.
5.5.3 Seed Selection Algorithm
Capitalising on Data Clustering and Graph Analysis, an algorithm (2) is developed for seed selection to achieve objectives (I) & (II). Given a target number of seeds to label, $T$, and the embeddings of all ESG concepts, $V$, seeds, $S_{total}$, are selected from clusters, $C$. Each cluster, $c_{k}$, within $C$, is a set of embeddings of concepts, $e_{i}$. Within each $c_{k}$, seeds of a cluster, $S_{c_{k}}$, are selected by taking the embedding with the maximum local consequential score $Q(e_{i})$ at each selection step. This maximises the selection of seeds with unique edges to increase label propagation to non-seeds. Additionally, each cluster should ideally possess an equivalent proportion, $P$, of seeds, and at least one seed each. We accomplish this by iteratively adjusting $P$ according to parameter $\beta$ (ESGSenticNet uses $\beta=0.01$), to ensure that our estimated number of seeds $N_{s}$ is as close to our target number as possible $T$ (the difference between $N_{s}$ and $T$ denoted by $\Delta$). From $P$, the specific number of seeds for each cluster, $n_{c_{k}}$ can be computed. This ensures that seeds are fairly represented across different clusters to enable the semantic diversity of seeds.
@@ 5.5.3 Seed Selection Algorithm
3:$N\leftarrow$ $|V|$
4:$P\leftarrow\frac{T}{N}$
5:$\Delta_{prev}\leftarrow\infty$

### v2

@@ 4.2 Taxonomy Construction Methodology
This approach first involves tokenising each sustainability report into separate sentences, with each sentence transformed into embeddings by Sentence-BERT [62]. BERTopic is then run on these embeddings, employing the KeyBERT configuration. The topic words are then provided to sustainability experts (details of the sustainability experts can be found in B) who refine the topic words into additional topics. Examples of topic words and their inferred topics can be found in Table 6:
Topic keywords Topic
@@
Table 6: Examples of Topics Inferred from Topic Keywords
Additional topics such as these are then added to an adapted version of the existing SASB framework [25], forming a taxonomy of ESG topics (Section 4.3) utilised in ESGSenticNet.
4.3 Resulting Taxonomy
Topic Type Topics
Pillar ‘Environmental’, ‘Social’, ‘Governance’.
Sub ‘Green Technologies’, ‘Green Infrastructure’, ‘Green Financing’, ‘Sustainable Corporate Operations’, ‘Sustainable Supply Chain’, ‘Sustainable Production Processes’, ‘Sustainable Land Management’, ‘Workplace Wellness’, ‘Workplace Diversity’, ‘Workplace Development’, ‘Community Empowerment’, ‘Strategic Partnerships’, ‘Customer Engagement’, ‘Management Composition’, ‘Management Ethics’, ‘Management Compensation’, ‘Stakeholder Engagement’, ‘Oversight’, ‘Environmental Compliance’, ‘Worker & Consumer Safety’, ‘Business Compliance’, ‘Data Privacy & Cybersecurity Protection’
Cross-Sub ‘Renewable and Efficient Energy’, ‘Material Sustainability’, ‘Water Conservation’, ‘Wastewater Management’, ‘Solid Waste Management’, ‘Climate Emissions’, ‘Air Quality Emissions’
Table 7: Topic types and topics, with the different topics delineated by a comma
We summarise the hierarchical taxonomy derived for ESGSenticNet in Table 7. The taxonomy comprises of topics are divided according to their different topic types–pillar, broad, cross-broad, sub, cross-sub, as shown in Table 7. The taxonomy is structured hierarchically, with topics categorised into different levels based on their scope and relationship. At the highest level, all topics fall under the broad ESG pillars—Environmental (E), Social (S), and Governance (G)—which represent the overarching dimensions of sustainability. Beneath the pillars, topics are grouped into broad and cross-broad topics, which define major sustainability themes. Each broad or cross-broad topic may have constituent sub-topics, which represent more specific areas within that theme. While Table 7 summarises our taxonomy, it does not detail the constituent sub-topics of each broad or cross-broad topic. For these relationships as well as topic definitions, please refer to the full descriptions in A and Table 16. We elaborate more on the features of this taxonomy below:
  • Realistic: Given that we leverage the empirical content from the sustainability reports, our taxonomy comprises additional ESG topics that might otherwise be overlooked by standard frameworks. For instance, ‘Green Technology’ is derived as a taxonomy topic, while not being explicitly emphasised in standard ESG reporting frameworks—i.e. GRI [43], SASB [25].
  • Activities Cross-Theme: To capture the interconnectedness of sustainability dimensions [48], additional topics include areas of a firm’s business activities (‘operations’, ‘development’ etc.). In contrast with existing frameworks that primarily describe direct sustainability impacts (‘emissions control’, ‘waste management’ etc.), this offers insight into the sustainability effects of specific business activities. In our taxonomy, topics pertaining to strategic business activities fall under ‘broad’ & ‘sub’ topic types, and are kept separate from topics pertaining to direct sustainability impacts, which fall under ‘cross-broad’ & ‘cross-sub’ topic types.
@@ 4.4 ESGSenticNet Relations & Rules
With the taxonomy established (Table 7), we now define the relation types between concept and pillar/topic (Table 8), as well as the rules governing these relations. Each concept within a triple can possess one of the following relations with a pillar/ topic:
@@
supports The concept directly advances the ESG topic with a clear and immediate positive impact.
undermines The concept directly impedes the ESG topic with a clear and immediate negative impact.
@@
The relations, ‘supports’ and ‘undermines’, clarify how concepts advance or impede crucial topics of sustainability. In contrast, the relation ‘aligns with’ does not consider a concept’s impact, highlighting a concept’s general connection with broader pillars of environmental, social or governance. At this juncture, it is worthwhile to qualify that the deployment of ESGSenticNet during topic analysis solely uses the ‘supports’ relation (explained further in Section 2.2), while the other relations serve as supplementary information. Separately, we now outline the rules that govern all relations in ESGSenticNet:
  • Pillar Assignment: Each concept ‘aligns with’ a specific pillar, mapping the concepts to the broader sustainability context.
  • Single Label within Topic Types: Each concept can hold only one relation with one topic within each topic type.
  • Cross-Labels between Parent & Children: Each concept can hold relations with higher-level topics (‘broad’ and ‘cross-broad’), as well as with more detailed topics (‘sub’ and ‘cross-sub’).
  • Cross-Labels between Cross & Non-Cross: Concepts that ‘align with’ the ‘environmental’ pillar can hold relations with ‘cross-broad’ and ‘cross-sub’ topics, in addition to ‘broad’ and ‘sub’ topics. This reflects the multi-faceted nature of environmental concepts, which can simultaneously impact environmental stewardship (represented by ‘cross-broad’, ‘cross-sub’ topics) and strategic business activities (represented by ‘broad’ and ‘sub’ topics).
4.5 ESGSenticNet Construction through a Neurosymbolic Framework
In sections 4.1 4.2 4.3 4.4, we outlined the preliminaries by introducing the data and defining the ESG topics and relations within the knowledge triples (concept, relation, ESG topic) in ESGSenticNet, thereby establishing ESGSenticNet’s organisational structure. In this section, we go beyond definitions to detail our neurosymbolic framework for deriving the triples, explaining how concepts are identified and assigned their respective relations.
Figure 6: Overview of neurosymbolic framework for ESGSenticNet construction
The neurosymbolic framework for constructing ESGSenticNet involves three main phases as shown in figure 6. Step 1: ESG Concept Parser (Section 4.6), involves concept parsing, which encompasses extracting candidate concepts from sustainability reports through dependency parsing and part-of-speech tagging. Step 2: Quality Control (Section 4.7), involves processing coherent concepts from these candidate concepts to ensure their coherence, as a form of quality control. Step 3: Labelling Process (Section 4.8), involves a labelling methodology to classify the filtered concepts according to the topics within our constructed taxonomy in Section 4.3, leveraging the rules defined in Section 4.4.
Accordingly, this labelling methodology in step (3) is detailed in Figure 6 (following processes i through iv). These processes integrate GPT-4o inference with a graph-based semi-supervised method, as explained below:
  1. (i)
    Semantic Graph Construction (Section 4.8.1): A semantic graph is constructed using the filtered concepts from step (2), leveraging their semantic similarities.
  2. (ii)
    Seed Selection (Section 4.8.2): Seed concepts are selected from the constructed semantic graph using a seed selection algorithm that applies clustering techniques and graph analysis.
  3. (iii)
    Seed Annotation (Section 4.8.3): Using GPT-4o, the relations between selected seed concepts and taxonomy topics are labelled.
  4. (iv)
    Label Propagation (Section 4.8.4): The relation labels from seed concepts are propagated to other concepts through a label propagation algorithm.
Upon completing the labelling process, ESGSenticNet is fully constructed. Notably, this labelling process allows us to label a large amount of diverse concepts despite performing inference on a subset of the concept samples. As a result, computational cost constraints are mitigated, as we are able to label concepts more extensively. This, in turn, significantly increases the extensiveness and diversity of our knowledge base, enhancing its effectiveness as a lexical method for ESG topic analysis.
@@ 4.6 ESG Concept Parser
Multiword expressions are linguistic phenomena that refer to common linguistic constructions such as idioms, collocations and sayings [19]. In the context of sustainability analysis, multiword expressions frequently communicate key sustainability ideas [65]. We refer to these multiword expressions as ESG “concepts”. Extracting these concepts or “concept extraction” has been explored for metaphors [46], finance [24] and sentiment analysis [13]. Yet, to the best of our knowledge, concept extraction within the ESG domain has not been studied. Therefore, to capture ESG concepts, a new concept parser is required, with an explicit focus on the distinct elements of sustainability discourse. These include the following designs:
@@
To effectively extract these unique aspects of sustainability language, we formalise ESG concepts through the following syntactic structure–verbs acting on complex noun phrases, where the words involved function as modifiers, descriptive and prescriptive elements. Specifically, the concepts are two or three word phrases, typically beginning with a verbverb, followed by $noun+noun$, or $adjective+noun$ combinations with specific grammatical dependencies. These dependencies include: verbnounverb-noun connection via (nsubj, obj, obl), adjectivenounadjective-noun or nounnounnoun-noun connection via (compound, amod, nn, appos, flat, nmod). This framework underpins the ESG Concept Parser algorithm (1), designed to capture key ideas within sustainability reports, particularly ESG efforts. For instance, ‘improve (VB) workplace (NN) safety (NN)’, where ‘workplace’ is the compoundcompound of ‘safety’ and ‘safety’ is the dobjdobj of ‘improve’.
@@
1:Input: Sentence from sustainability report
@@
[…]
@@
[…]
@@
4.7 Quality Control
To minimise computational cost, parsed concepts from algorithm (1) are filtered to retain the top 110k most frequently occurring ones. Of these, we eventually annotate 23k concepts based on our estimated computational resources, which dictates the number of seeds we select within our semantic graph. Prior to the annotation process, these concepts must undergo processing to ensure their coherence, as a form of quality control. This involves utilising GPT-4o in few-shot setting (Brown, 2020). GPT-4o is leveraged due to the capacity of large generative LLMs to infer [72], as well as their powerful natural language understanding capabilities [76]. Using GPT-4o to process concepts for coherence involves conducting the following inference tasks in order: (1) reorder words, if required, to improve the coherence of the concept, (2) determine if the subsequent concept is intelligible. We provide the prompt for (1) and (2) through Figure 7 and Figure 8 respectively.
@@ 4.7 Quality Control
4.8 Labelling Process
4.8.1 Semantic Graph Construction
@@
$w_{ij}=\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j}),\quad\forall(\mathbf{e}_{i},\mathbf{e}_{j})\in E$ (5)
Processed concepts undergo labelling to derive their relationships with different ESG topics. Given the exceedingly large number of concepts, annotating all concepts would be prohibitively expensive. To overcome this, we devise a semi-supervised framework that enables a large volume of diverse concepts to be labelled from annotating only a subset of all concepts. Via a semantic graph, we leverage the semantic similarities between concepts to propagate labels from annotated concepts (seeds) to unlabelled concepts (non-seeds). This graph is constructed by transforming each concept $cp_{i}$ into embeddings $e_{i}$ through S-BERT [62], with edges constructed by leveraging a cosine similarity threshold of 0.80 between each node $e_{i}$. To qualify the strength of relation between nodes (concepts), the edge weights $w_{ij}$ are defined as the cosine similarity scores between the nodes. We describe the full graph via equations (1) through (5).
4.8.2 Seed Selection
Within our semantic graph, seeds are selected for annotation through data clustering and graph analysis. Given the semi-supervised label propagation from seeds to non-seeds, we formulate the following objectives for seed selection: (I) Seeds must be semantically diverse to ensure that diverse labelled samples are yielded post label propagation. This enhances the variety and comprehensiveness of labelled concepts in our knowledge base. (II) Seeds must propagate their labels to as many non-seeds as possible. This maximises the number of labelled samples we can attain while minimising annotation costs. To achieve (I) and (II), we utilise Data Clustering and Graph Analysis, with these components culminating in a Seed Selection Algorithm.
@@ 4.8.2 Seed Selection
Data Clustering. To accomplish (I), the embeddings of concepts $e_{i}$ are clustered to enable seed selection from diverse semantic spaces. This clustering involves projecting the embeddings to a lower dimensional space via UMAP [47], before running HDBSCAN [15] on these projected embeddings. While HDBSCAN is flexible to the shape and number of clusters, and robust to noise [15], reducing the dimensionality of embeddings ensures good clustering performance. This is because HDBSCAN relies on distance-based metrics which lack meaningfulness in the high dimensional spaces of textual embeddings [3]. The embeddings are projected to lower dimensions for the sole purpose of clustering. In our later section, label propagation leverages the default dimensions of S-BERT embeddings. Clusters are evaluated via a Confidence Quality Index ($CQI$), which denotes the proportion of embeddings, $e$, out of all embeddings, $V$, with a clustering confidence level $\alpha_{e}$ beyond 0.60, as described by equation (6). For our use of HDBSCAN, this simple metric is more appropriate than others like Slihouette Score [63], which assumes that all embeddings belong to a cluster. Hyperparameter tuning is done via random search, with the best run having a $CQI$ of 0.66, UMAP parameters nn_neighboursneighbours==32, nn_componentscomponents==2, and the HDBSCAN parameter minmin_clustercluster_sizesize==2.
@@
Graph Analysis. To accomplish (II), graph analysis is utilised to select seeds that are the most consequential nodes. Given that edges facilitate label propagation, a node is considered more consequential if it possesses a greater number of unique edges. Unique edges are defined as connections to nodes that do not share an edge with the already selected seed nodes within the same cluster. We formulate the local consequential score $Q(e_{i})$, where $S_{c_{k}}$ is the set of nodes already selected as seeds within the same cluster, $c_{k}$, as the node $e_{i}$, with this described in equation (7).
Seed Selection Algorithm. Capitalising on Data Clustering and Graph Analysis, an algorithm (2) is developed for seed selection to achieve objectives (I) & (II). Given a target number of seeds to label, $T$, and the embeddings of all ESG concepts, $V$, seeds, $S_{total}$, are selected from clusters, $C$. Each cluster, $c_{k}$, within $C$, is a set of embeddings of concepts, $e_{i}$. Within each $c_{k}$, seeds of a cluster, $S_{c_{k}}$, are selected by taking the embedding with the maximum local consequential score $Q(e_{i})$ at each selection step. This maximises the selection of seeds with unique edges to increase label propagation to non-seeds. Additionally, each cluster should ideally possess an equivalent proportion, $P$, of seeds, and at least one seed each. We accomplish this by iteratively adjusting $P$ according to parameter $\beta$ (ESGSenticNet uses $\beta=0.01$), to ensure that our estimated number of seeds $N_{s}$ is as close to our target number as possible $T$ (the difference between $N_{s}$ and $T$ denoted by $\Delta$). From $P$, the specific number of seeds for each cluster, $n_{c_{k}}$ can be computed. This ensures that seeds are fairly represented across different clusters to enable the semantic diversity of seeds.
@@
[…]
@@
4.8.3 Seed Annotation
Figure 9: Flowchart depicting inference process
Through GPT-4o few-shot prompting, the selected seeds are annotated with their relations (‘aligns with’, ‘supports’, ‘undermines’) with respect to a pillar or topic, before their labels are later propagated within the semantic graph (Section 4.8.4). ESGSenticNet retains a considerable number of seed concepts (>16k>16k) directly labelled through GPT-4o. This improves the accuracy of ESGSenticNet’s relations, given that the reasoning capabilities of GPT-4o [36] allows it to infer relations for concepts outside of similar semantic spaces. For example, consider the concepts ‘use led light’, ‘achieve energy efficiency’ categorised under ‘Resource Optimisation’. Textual similarity methods leveraging these for comparison might not categorise ‘emphasise water importance’ similarly, despite its relevance. However, using GPT-4o, ‘emphasise water importance’ is appropriately labelled, due to the LLM’s capacity to infer semantic relationships beyond mere textual similarities.
To improve GPT-4o classification accuracy, inference tasks are divided into smaller, more manageable sub-tasks [17]. These tasks align with the taxonomy’s hierarchy levels, following the flowchart in Figure 9. Subsequent categorisations of seeds are contingent upon their initial classification into broader topics, and it aligns with the predefined associations between the ESG pillars and broad topics, as well as the connections between the broad and sub topics. Prompts for these tasks are highlighted in Figure 10 (pillar labelling), Figure 11 (relation-topic labelling).
[Task Description]
As a sustainability expert, your task is to classify phrases into the topics of Environmental, Social, Governance, or Others based on their explicit relevance to the Environmental, Social, Governance topics.
[Definitions of topics]
Others: Reserved for phrases that do not meet the explicitness and specificity requirements for Environmental, Social, or Governance. This includes phrases that are vague, ambiguous, or cover multiple Environmental, Social, Governance topics without clear categorisation.
[Response Instructions]
Evaluate each phrase’s explicit content to determine its most appropriate topic. Assign each phrase to only one topic: Environmental, Social, Governance, or Others. Only phrases that are clear and explicit can be categorised under Environmental, Social, or Governance. Phrases that are vague or not clearly related to these topics should be categorised as Others. Ensure all responses strictly adhere to the definitions and format provided.
[Response Format]
Input: <phrase>
Response: <Environmental, Social, Governance, or Others>
Figure 10: Prompt for labelling an ESG concept’s pillar
[Task Description] As a sustainability expert, you are tasked with evaluating the relationship between phrases and ESG-related topics. [Definitions of Relationships] supports: The phrase, specific and explicit to the topic, directly advances it with a clear and immediate positive impact. undermines: The phrase, specific and explicit to the topic, directly impedes its goals with a clear and immediate negative impact. [Definitions of Topics] Definitions of topics according to taxonomy [Response Instructions] Evaluate the explicit content of each phrase to determine the most appropriate and accurate relationship and topic. Identify the single most relevant and appropriate relationship and topic. If no specific and explicit relationship can be determined between the phrase and any topic, or if the phrase does not fit any topic due to vagueness or irrelevance, classify the phrase as ‘not applicable’. [Response Format] For each phrase, provide your analysis using the format below. Output only one tuple if applicable, otherwise state ‘not applicable’. Input: <phrase> Response: <(relationship, topic)> or ‘not applicable’
Figure 11: Prompt for labeling an ESG concept’s relation with a topic
4.8.4 Label Propagation
$L^{(k+1)}=D^{-1/2}AD^{-1/2}L^{(k)},$ (8)
$l_{i}^{(k+1)}=l_{i}^{(0)}\quad\forall i\leq m$ (9)
Annotated seeds have their labels propagated to non-seeds via the label propagation algorithm [69]. This increases the labelled data within our knowledge base without incurring additional labelling costs. We utilise the label propagation algorithm for its ability to perform soft transducive inference — iteratively proliferating labels across the graph probabilistically, based on global similarity [69]. This contrasts with other algorithms — i.e. k-nearest neighbours [59], which assign hard labels without modeling uncertainty (i.e. via probability). In our context, the label propagation algorithm’s incorporation of probability allows it to better capture semantic nuances and ambiguity across concepts. This is useful when one concept may relate to multiple categories. Specifically, the following algorithm is iterated until convergence, where $A$ is an adjacency matrix containing $a_{i,j}$ that denotes the edge weight between connected nodes $e_{i}$ and $e_{j}$. Each node represents an embedding of a concept, while the edge weight denotes cosine similarity between concepts. $L$ is a matrix that represents the relations labels of each node with respect to a topic (i.e. supports within the (‘reduce water consumption’, supports, ‘resource optimisation’). $m$ is the number of nodes possessing labels and $D$ is a diagonal matrix containing $d_{i,i}$, which equals the sum of elements in row $i$ of $A$. We use parameters n_layersn\_layers=50, $\alpha$=0.5.
4.9 ESGSenticNet — Statistics & Samples
Figure 12: Overview of ESGSenticNet
After label propagation, ESGSenticNet is fully constructed. We provide an visualisation of ESGSenticNet in Figure 12. ESGSenticNet comprises 44232 triples containing 23245 unique ESG concepts. Of the 23245 concepts, 16011 are seeds and 7234 are non-seeds. We provide samples of ESGSenticNet’s knowledge triples (concept, relation, ESG topic) in Table 9.
Concept Relation ESG topic
build diverse workplace supports workplace diversity
halve carbon emission supports emissions control
minimise resource consumption supports resource optimisation
organise charity event supports outreach
produce chemical waste undermines operations
Table 9: Samples of knowledge triples from ESGSenticNet
Table 10 provides a meta statistic of of how ESGSenticNet compares with existing sustainability dictionaries—i.e Baier [5], Naiara [60], Kang [39]. To qualify, the concepts within ESGSenticNet’s triples are taken as “terms”.
Sustainability Dictionaries/ Knowledge Base(s) Total Unique Terms
ESGSenticNet 23245 (within 44232 triples)
Baier 491
Naiara 167
Kang* 792
Table 10: Total number of terms of different sustainability corpora, *indicates not publicly available.
Finally, we provide a full statistical breakdown of ESGSenticNet in Appendix C Table 17. This outlines the total triples corresponding to each ESG topic, with a breakdown of the relations involved.
4.10 Ablation Studies